Ratio of the pressures of the two gases if 10 g of each methane and oxygen are kept in cylinders having same volume and have same values of temperature?
1
Expert's answer
2019-04-17T04:21:34-0400

PV=nRT;

V = const, R = const, T = const;

If 1 – is methane (CH4), and 2 – oxygen (02), so

P1*V=n1*R*T;

P2*V=n2*R*T;

P1/P2 = n1/n2 = (m1/M1)/(m2/M2);

m1 = m2 = 10 g, so

P1/P2 = M2/M1;

M1 (methane (CH4)) = 16 g/mol and M2 (oxygen (02)) = 32 g/mol;

P1/P2 = M2/M1 = 32/16 = 2.

Answer: 2.